### v1

@@ B.4 Simulations
where $p_{d}$ is the dark count probability of Bob’s detectors, $e_{\rm mis}$ is the misalignment error rate of the overall system and $\eta$ is the overall system loss. Namely, $\eta=10^{-\lambda/10}$, with $\lambda$ denoting the loss in the channel in dB.
This model is used to select the threshold values of the PE test. To be precise, these thresholds are set to the expectation values of the underlying variables according to the channel model. As an example, within the PE test $\Omega_{\rm TEST}=\{n_{1,\rm Z}^{\rm L}\geq n_{1,\rm Z}^{\rm th},\ \phi_{1,\rm Z}^{\rm U}\leq\phi_{1,\rm Z}^{\rm th}\}$, $n_{1,\rm Z}^{\rm th}$ is set to the decoy-state bound of Eq. (14), upon replacement of $n_{{\rm Z},k}$ by its expectation according to the channel model:
$\mathbb{E}\left[n_{{\rm Z},k}\right]=N_{\rm Z}\frac{p_{k}D_{k}}{\sum_{j}p_{j}D_{j}}.$ (40)
Similarly, $\phi_{1,\rm Z}^{\rm th}$ is set to the threshold function derived in this work, upon replacement of all the observables ($n_{{\rm Z},k}$, $n_{{\rm X},k}$ and $m_{{\rm X},k}$) by their expectations according to the channel model. Particularly,
$\mathbb{E}\left[m_{{\rm X},k}\right]=N_{\rm X}\frac{p_{k}e_{k}}{\sum_{j}p_{j}D_{j}},$ (41)
and $\mathbb{E}[n_{{\rm X},k}]$ is obtained simply replacing $Z$ by $X$ in Eq. (40).
We remark that the above choices of the thresholds in the PE test are over-optimistic, as they would incur in large abortion probabilities even if the channel behaves according to the considered model. In fact, the “robustness” penalty of using these thresholds is certainly larger when considering the more convoluted PE test devised for the random sampling tool of [22]. This is so because, in that test, double-sided thresholds are imposed, and since both the upper and the lower thresholds are being set to the same value for the simulations, the success probability of the test becomes negligible. In either case, a more sensitive approach consists of making robustness considerations to incorporate statistical fluctuations in the thresholds (ultimately, however, in a real experiment, the thresholds would be chosen via careful channel monitoring).
This said, in Fig. 4 below we plot the secret key rate as a function of the block length $N=N_{\rm Z}+N_{\rm X}$ for the considered decoy-state BB84 protocol. When compared to Fig. 2b of the main text, this figure introduces two additional lines, where we combine the Neyman constructions of Proposition 1 of the main text with either Ref. [22] or [15] for the random sampling bound. These new lines highlight the advantage of our random sampling tool when compared to Ref. [22] or [15] for decoy-state QKD.
Figure 4: Finite secret key rate per raw data bit as a function of the data block size for the decoy-state BB84 protocol. The solid blue line corresponds to Propositions 1 and 2 of the main text, the dashed orange (dotted green) line combines [22] with [30] ([24]), while the red (purple) line combines Proposition 1 with [22] ([15]). The experimental parameters considered in the simulations are the same as in the main text.
Appendix C Neyman constructions for the additive Chernoff bounds
In the first place, we provide a relaxation of the upper additive Chernoff bound [26].
@@ Appendix C Neyman constructions for the additive Chernoff bounds
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, such that $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $z\in[p,1]$, $\Pr[\hat{p}\geq{}z]\leq{e^{-n{}D(z,p)}}$ for $D(z,p)=9(z-p)^{2}/2(z+2p)(3-z-2p)$.
@@
Using the logarithmic inequality $\ln(1+z)\geq z(6+5z)/2(1+z)(3+z)$ [27] within Eq. (42) yields $D(z||p)\geq{}9(z-p)^{2}/\left(2(z+2p)(3-z-2p)\right)=:D(z,p)$. Hence, the desired relaxation follows. ∎
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, such that $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $\epsilon>0$, $\Pr[\hat{p}\geq{}z_{n,\epsilon}(p)]\leq{\epsilon}$ with
$z_{n,\epsilon}(p)=\left\{\begin{array}[]{ll}\displaystyle{\frac{-b_{n,\epsilon}(p)+\sqrt{b_{n,\epsilon}(p)^{2}-4a_{n,\epsilon}c_{n,\epsilon}(p)}}{2a_{n,\epsilon}}}&\mathrm{if}\hskip 5.69046ptp\leq{}\displaystyle{\frac{1-2\kappa_{n,\epsilon}}{1+4\kappa_{n,\epsilon}}},\\ 1&\rm{otherwise,}\\ \end{array}\right.$ (43)
where $a_{n,\epsilon}=1+\kappa_{n,\epsilon}$, $b_{n,\epsilon}(p)=-2p-\kappa_{n,\epsilon}(3-4p)$ and $c_{n,\epsilon}(p)=p^{2}-\kappa_{n,\epsilon}{}p(6-4p)$ for $\kappa_{n,\epsilon}=(2/9n)\ln{\epsilon^{-1}}$ ).
@@
According to Proposition 1 above, for all $z\in[p,1]$, $\Pr[\hat{p}\geq{}z]\leq{\epsilon}$ holds if $\epsilon=\exp\{-n{}D(z,p)\}$. Let us consider the function $g_{n,p}(z)=\exp\{-n{}D(z,p)\}$ that maps $z$’s to $\epsilon$’s for abitrary $n$ and $p$. Since $g_{n,p}(z)$ is decreasing and thus injective in $[p,1]$, it is invertible in $g_{n,p}\left([p,1]\right)=\bigl{[}g_{n,p}(1),1\bigr{]}$. Thus, in virtue of Proposition 1, for all $\epsilon\in{}\bigl{[}g_{n,p}(1),1\bigr{]}$ $\Pr[\hat{p}\geq{}z]\leq{\epsilon}$ holds if $z=g^{-1}_{n,p}(\epsilon)$, which is the only $z\in[p,1]$ such that $g_{n,p}(z)=\epsilon$. This is a quadratic equation $a_{n,\epsilon}z^{2}+b_{n,\epsilon}(p){}z+c_{n,\epsilon}(p)=0$ with the coefficients prescribed in the statement, and one can readily show that only the larger root, say $z_{n,\epsilon}^{+}(p)$, lies in $[p,1]$. In short, $\Pr[\hat{p}\geq{}z_{n,\epsilon}^{+}(p)]\leq{\epsilon}$ for all $\epsilon\geq{}g_{n,p}(1)$. Lastly, the claim follows by noticing that $\epsilon\geq{}g_{n,p}(1)\Leftrightarrow{}p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$ and completing the missing range with the trivial bound [38]. ∎
Proposition 2 above can be used to infer a Neyman [25] construction on the underlying parameter $p$ in terms of $\hat{p}$. For this purpose, we introduce the function $\tilde{z}_{n,\epsilon}(p)$, which results from extending ${z}_{n,\epsilon}(p)$ with its last slope at the two points where it intersects the $[0,1]\times{}[0,1]$ square. Precisely, analytical inspection shows that ${z}_{n,\epsilon}(p)$ is monotonically increasing in the interval $[0,(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})]$, whose extreme points provide the intersections with the $[0,1]\times{}[0,1]$ square: $z_{n,\epsilon}(0)=3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$ and $z_{n,\epsilon}((1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon}))=1$. The motivation of “removing the kinks” like this is that Proposition 2 trivially follows as well if we replace $z_{n,\epsilon}$ by $\tilde{z}_{n,\epsilon}$, but $\tilde{z}_{n,\epsilon}$ has the advantage of being globally invertible, thus enabling the well-definiteness of the random variable $\Gamma^{-}_{n,\epsilon}(\hat{p})=\tilde{z}^{-1}_{n,\epsilon}(\hat{p})$ for all $\hat{p}\in[0,1]$.
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, such that $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p]\leq{\epsilon}$ with $\Gamma^{-}_{n,\epsilon}(\hat{p})=\tilde{z}^{-1}_{n,\epsilon}(\hat{p})$.
@@
Since $\tilde{z}^{-1}_{n,\epsilon}$ is monotonically increasing, it follows that $\left\{\tilde{z}^{-1}_{n,\epsilon}(\hat{p})\geq{}p\right\}\implies{}\left\{\hat{p}\geq{}\tilde{z}_{n,\epsilon}(p)\right\}$ irrespectively of $p$. Therefore, $\Pr\left[\tilde{z}^{-1}_{n,\epsilon}(\hat{p})\geq{}p\right]\leq{}\Pr\left[\hat{p}\geq{}\tilde{z}_{n,\epsilon}(p)\right]\leq{}\epsilon$, where in the last inequality we invoked the relaxation of Proposition 2 above that results from replacing $z_{n,\epsilon}$ by $\tilde{z}_{n,\epsilon}$. ∎
Let us now provide an explicit formula for the “principal branch” of $\Gamma^{-}_{n,\epsilon}(\hat{p})=\tilde{z}^{-1}_{n,\epsilon}(\hat{p})$, corresponding to the interval $\hat{p}\geq{}\tilde{z}_{n,\epsilon}(0)=3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$. Since $\tilde{z}^{-1}_{n,\epsilon}(\hat{p})$ is the only $p<\hat{p}$ such that $\tilde{z}_{n,\epsilon}(p)=\hat{p}$, solving this quadratic equation for $p$ (and choosing the relevant root) we conclude that
$\Gamma^{-}_{n,\epsilon}(\hat{p})=\frac{1}{1+4\kappa_{n,\epsilon}}\left(3\kappa_{n,\epsilon}+(1-2\kappa_{n,\epsilon})\hat{p}-3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+\hat{p}-\hat{p}^{2}\right)}\right),$ (44)
for $\hat{p}\geq{}3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$. For $\hat{p}<3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$, the function $\Gamma^{-}_{n,\epsilon}(\hat{p})$ is extended with the slope at $\hat{p}=3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon})$.
We remark that all the above propositions originate from the upper additive Chernoff bound. Naturally, for each of these propositions, there exists an analogous statement arising from the lower additive Chernoff bound instead. In particular, a Neyman construction complementary to that of Proposition 3 follows.
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, such that $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where
$\Gamma^{+}_{n,\epsilon}(\hat{p})=\frac{1}{1+4\kappa_{n,\epsilon}}\left(3\kappa_{n,\epsilon}+(1-2\kappa_{n,\epsilon})\hat{p}+3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+\hat{p}-\hat{p}^{2}\right)}\right),$ (45)
for $\hat{p}\leq{}(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$, and for $\hat{p}>(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$ $\Gamma^{+}_{n,\epsilon}(\hat{p})$ is extended with the slope at $\hat{p}=(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})$.
@@
The proof proceeds identically as that of Proposition 3 above, but using the lower additive Chernoff bound as the starting point. ∎
Propositions 3 and 4 above correspond to Proposition 1 in the main text.

### v2

@@ II.4 Channel model
where $p_{d}$ is the dark count probability of Bob’s detectors, $e_{\rm mis}$ is the misalignment error rate of the system, and $\eta$ is the overall system efficiency. Namely, $\eta=10^{-\lambda/10}$, $\lambda$ denoting the total loss in dB.
This model is used to select the threshold values of the PE test. To be precise, the thresholds are set to the expectation values of the underlying variables according to the channel model. As an example, within the PE test $\Omega_{\rm TEST}=\{n_{1,\rm Z}^{\rm L}\geq n_{1,\rm Z}^{\rm th},\ \phi_{1,\rm Z}^{\rm U}\leq\phi_{1,\rm Z}^{\rm th}\}$, $n_{1,\rm Z}^{\rm th}$ is set to the decoy-state bound of Eq. (19), upon replacement of $n_{{\rm Z},k}$ by its expectation according to the channel model:
$E\left[n_{{\rm Z},k}\right]=N_{\rm Z}\frac{p_{k}D_{k}}{\sum_{j}p_{j}D_{j}}.$ (47)
Similarly, $\phi_{1,\rm Z}^{\rm th}$ is set to the threshold function under consideration, upon replacement of all the observables ($n_{{\rm Z},k}$, $n_{{\rm X},k}$ and $m_{{\rm X},k}$) by their expectations according to the channel model. Particularly,
$E\left[m_{{\rm X},k}\right]=N_{\rm X}\frac{p_{k}e_{k}}{\sum_{j}p_{j}D_{j}},$ (48)
and $E[n_{{\rm X},k}]$ is obtained simply replacing $Z$ by $X$ in Eq. (47).
We remark that the above choices of the thresholds in the PE test are over-optimistic, as they would incur in large abortion probabilities even if the channel behaves according to the considered model [49]. A more sensitive approach consists of making robustness considerations to incorporate statistical fluctuations in the thresholds. Ultimately, however, in a real experiment, the thresholds would be chosen via careful channel monitoring.
Appendix III: Computing confidence intervals from the additive Chernoff bounds
This section includes the technical derivations underlying Proposition 1 in the main text. In the first place, we provide a relaxation of the upper additive Chernoff bound [20].
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $z\in[p,1]$, $\Pr[\hat{p}\geq{}z]\leq{e^{-n{}D(z,p)}}$ for $D(z,p)=9(z-p)^{2}/2(z+2p)(3-z-2p)$.
@@
Plugging the logarithmic inequality $\ln(1+z)\geq z(6+5z)/2(1+z)(3+z)$ [24] in each term of the divergence yields
$D(z||p)\geq z\frac{(\frac{z}{p}-1)(6+5(\frac{z}{p}-1))}{2(1+(\frac{z}{p}-1))(3+(\frac{z}{p}-1))}+(1-z)\frac{(\frac{1-z}{1-p}-1)(6+5(\frac{1-z}{1-p}-1))}{2(1+(\frac{1-z}{1-p}-1))(3+(\frac{1-z}{1-p}-1))},$ (50)
which, after some straightforward manipulation, leads to
$D(z||p)\geq{}\frac{(z-p)(p+5z)}{2(z+2p)}+\frac{(p-z)(6-5z-p)}{2(3-z-2p)}.$ (51)
Lastly, combining both summands in a single fraction and simplifying yields
$D(z||p)\geq\frac{9(z-p)^{2}}{2(z+2p)(3-z-2p)}=:D(z,p)$ (52)
Hence, the desired relaxation follows. ∎
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left(\hat{p}\right)=p$ such that $p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$ for $\kappa_{n,\epsilon}=(2/9n)\ln(1/\epsilon)$. Then, $\Pr[\hat{p}\geq{}z_{n,\epsilon}(p)]\leq{\epsilon}$ with
$z_{n,\epsilon}(p)=\displaystyle{\frac{-b_{n,\epsilon}(p)+\sqrt{b_{n,\epsilon}(p)^{2}-4a_{n,\epsilon}c_{n,\epsilon}(p)}}{2a_{n,\epsilon}}},$ (53)
where $a_{n,\epsilon}=1+\kappa_{n,\epsilon}$, $b_{n,\epsilon}(p)=-2p-\kappa_{n,\epsilon}(3-4p)$ and $c_{n,\epsilon}(p)=p^{2}-\kappa_{n,\epsilon}{}p(6-4p)$.
@@
According to Proposition 1, for all $z\in[p,1]$, $\Pr[\hat{p}\geq{}z]\leq{\epsilon}$ holds if $\epsilon=\exp\{-n{}D(z,p)\}$. Let us consider the function $g_{n,p}(z)=\exp\{-n{}D(z,p)\}$ that maps $z$’s to $\epsilon$’s for arbitrary $n$ and $p$. Since $g_{n,p}(z)$ is decreasing and thus injective in $[p,1]$, it admits an inverse in $g_{n,p}\left([p,1]\right)=\bigl{[}g_{n,p}(1),1\bigr{]}$. Namely, for all $\epsilon\geq{}g_{n,p}(1)$, $\Pr[\hat{p}\geq{}z]\leq{\epsilon}$ holds if $z=g^{-1}_{n,p}(\epsilon)$, which is the only $z\in[p,1]$ such that $g_{n,p}(z)=\epsilon$. This is a quadratic equation $a_{n,\epsilon}z^{2}+b_{n,\epsilon}(p){}z+c_{n,\epsilon}(p)=0$ with the coefficients prescribed in the statement, and one can readily show that only the larger root lies in $[p,1]$. Lastly, the claim follows by observing that $\epsilon\geq{}g_{n,p}(1)\Leftrightarrow{}p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$ [50]. ∎
Let us now use Proposition 2 to infer a one-sided CI for the underlying parameter $p$ [51, 52]. Analytical inspection shows that ${z}_{n,\epsilon}(p)$ is monotonically increasing in the interval $[0,(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})]$, and thus it admits an inverse $z^{-1}_{n,\epsilon}$ in $I_{n,\epsilon}={z}_{n,\epsilon}([0,(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})])=\left[3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon}),1\right]$. By definition, $z^{-1}_{n,\epsilon}(x)$ is the only $p<x$ such that $z_{n,\epsilon}(p)=x$. Solving this quadratic equation for $p$ and choosing the relevant root we conclude that, for all $x\in\left[3\kappa_{n,\epsilon}/(1+\kappa_{n,\epsilon}),1\right]$,
$z^{-1}_{n,\epsilon}(x)=\frac{1}{1+4\kappa_{n,\epsilon}}\left(3\kappa_{n,\epsilon}+(1-2\kappa_{n,\epsilon})x-3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+x-x^{2}\right)}\right).$ (54)
Essentially, $z^{-1}_{n,\epsilon}(\hat{p})$ is the statistic we use to construct a CI on $p$, as shown next.
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left[\hat{p}\right]=p$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p]\leq{\epsilon}$, where $\Gamma^{-}_{n,\epsilon}(\hat{p})=z^{-1}_{n,\epsilon}(\hat{p})$ if $\hat{p}\in{}I_{n,\epsilon}$, and $\Gamma^{-}_{n,\epsilon}(\hat{p})=-\epsilon$ otherwise.
@@
For all $\hat{p}$, $\Gamma^{-}_{n,\epsilon}(\hat{p})\leq{}\Gamma^{-}_{n,\epsilon}(1)=(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$. Hence, the claim holds if $(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})<p\leq{}1$. Let us now consider $0\leq{}p\leq{}(1-2\kappa_{n,\epsilon})/(1+4\kappa_{n,\epsilon})$. We have $\Pr\left[\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p\right]=\Pr\left[\hat{p}\in{}I_{n,\epsilon},\Gamma^{-}_{n,\epsilon}(\hat{p})\geq{}p\right]=\Pr\left[\hat{p}\in{}I_{n,\epsilon},z^{-1}_{n,\epsilon}(\hat{p})\geq{}p\right]\leq{}\Pr\left[\hat{p}\in{}I_{n,\epsilon},\hat{p}\geq{}z_{n,\epsilon}(p)\right]\leq{}\Pr\left[\hat{p}\geq{}z_{n,\epsilon}(p)\right]\leq{}\epsilon.$ Here, the first equality follows because $\hat{p}\notin{}I_{n,\epsilon}\implies{}\Gamma^{-}_{n,\epsilon}(\hat{p})=-\epsilon<p$ [53], the second equality follows because $\hat{p}\in{}I_{n,\epsilon}\implies\Gamma^{-}_{n,\epsilon}(\hat{p})=z^{-1}_{n,\epsilon}(\hat{p})$, the first inequality follows because $\{\hat{p}\in{}I_{n,\epsilon},z^{-1}_{n,\epsilon}(\hat{p})\geq{}p\}\implies{}\{\hat{p}\in{}I_{n,\epsilon},\hat{p}\geq{}z_{n,\epsilon}(p)\}$ —as $z^{-1}_{n,\epsilon}(\hat{p})$ is monotonically increasing in $I_{n,\epsilon}$—, the second inequality follows because $\{\hat{p}\in{}I_{n,\epsilon},\hat{p}\geq{}z_{n,\epsilon}(p)\}\implies{}\{\hat{p}\geq{}z_{n,\epsilon}(p)\}$, and the final inequality follows from Proposition 2 above. ∎
We remark that, ultimately, propositions 1 to 3 originate from the upper additive Chernoff bound. Importantly, for each of them, there exists an analogous statement arising from the lower additive Chernoff bound instead. In particular, a one-sided CI complementary to that of Proposition 3 follows.
@@
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left(\hat{p}\right)=p$. Then, for all $\epsilon>0$, $\Pr[\Gamma^{+}_{n,\epsilon}(\hat{p})\leq{}p]\leq{\epsilon}$, where
$\Gamma^{+}_{n,\epsilon}(\hat{p})=\frac{1}{1+4\kappa_{n,\epsilon}}\left(3\kappa_{n,\epsilon}+(1-2\kappa_{n,\epsilon})\hat{p}+3\sqrt{\kappa_{n,\epsilon}\left(\kappa_{n,\epsilon}+\hat{p}-\hat{p}^{2}\right)}\right)$ (55)
if $\hat{p}\in[0,(1-2\kappa_{n,\epsilon})/(1+\kappa_{n,\epsilon})]$, and $\Gamma^{+}_{n,\epsilon}(\hat{p})=1+\epsilon$ otherwise.
@@
The proof follows similar lines as that of Proposition 3 above, but using the lower additive Chernoff bound as the starting point. ∎
Proposition 1 in the main text is the conjunction of propositions 3 and 4 presented here.
Appendix IV: Optimal confidence intervals for the average parameter of a set of independent Bernoulli variables
Here, we present the tightest monotonic one-sided CIs for the average parameter of a set of independent Bernoulli variables [54, 55, 56], originally derived in [30].
Theorem 1.
Let $\hat{p}$ be the average of $n$ independent Bernoulli variables, with expected value $\mathbb{E}\left[\hat{p}\right]=p$. Then, for all $\epsilon\in(0,1/4]$, we have that $\Pr[\mathcal{F}^{-}_{n,\epsilon}(\hat{p})>p]\leq{\epsilon}$ for
$\mathcal{F}^{-}_{n,\epsilon}(\hat{p})=\left\{\begin{array}[]{ll}0&\mathrm{if}\hskip 5.69046ptn\hat{p}=0,\\ \hat{p}-\displaystyle{\frac{1-\epsilon}{n\left[1-\epsilon^{*}(n\hat{p},n)\right]}}&\mathrm{if}\hskip 5.69046ptn\hat{p}>0,\hskip 5.69046pt\epsilon^{*}(n\hat{p},n)\leq{}\epsilon\leq{}1,\\ I^{-1}_{\epsilon}(n\hat{p},n-n\hat{p}+1)&\mathrm{if}\hskip 5.69046ptn\hat{p}>0,\hskip 5.69046pt0\leq{}\epsilon\leq{}\epsilon^{*}(n\hat{p},n),\end{array}\right.$ (56)
where $\epsilon^{*}(n\hat{p},n)=I_{(n\hat{p}-1)/n}(n\hat{p},n-n\hat{p}+1)$ and $I^{-1}_{\epsilon}(a,b)$ denotes the inverse regularized incomplete beta function, such that $I_{I^{-1}_{\epsilon}(a,b)}(a,b)=\epsilon$ for
[…]
Complementarily, for all $\epsilon\in(0,1/4]$, we have that $\Pr[\mathcal{F}^{+}_{n,\epsilon}(\hat{p})<p]\leq{\epsilon}$ for $\mathcal{F}^{+}_{n,\epsilon}(\hat{p})=1-\mathcal{F}^{-}_{n,\epsilon}(1-\hat{p})$.
To finish with, two observations are in order. On the one hand, note that for any $\Delta>0$, $\left\{\mathcal{F}^{-}_{n,\epsilon}(\hat{p})\geq{}p+\Delta\right\}\implies{}\left\{\mathcal{F}^{-}_{n,\epsilon}(\hat{p})>p\right\}$, such that $\Pr[\mathcal{F}^{-}_{n,\epsilon}(\hat{p})\geq{}p+\Delta]=\Pr[\left(\mathcal{F}^{-}_{n,\epsilon}(\hat{p})-\Delta\right)\geq{}p]\leq{\epsilon}$ in virtue of Theorem 1. In other words, by decreasing $\mathcal{F}^{-}_{n,\epsilon}(\hat{p})$ in any nonzero amount $\Delta$, we reach a looser lower bound that trivially sticks to the sign convention used in this work, i.e. contemplating the equality sign in the probability (rather than the strict inequality considered in Theorem 1). Similarly, increasing $\mathcal{F}^{+}_{n,\epsilon}(\hat{p})$ in any nonzero amount allows to meet the desired criterion for the upper bound.
On another note, binomial-like CIs slightly simpler than those of Theorem 1 are provided in [57], which only entail a trivial loss of optimality imperceptible with the sample sizes characteristic of QKD applications.